Hahn Company uses a job-order costing system. Its plantwide predetermined overhead rate uses direct labor-hours as the allocation base. The company pays its direct laborers $16.50 per hour. During the year, the company started and completed only two jobs-Job Alpha, which used 61,700 direct labor-hours, and Job Omega. The job cost sheets for the these two jobs are shown below: Job Alpha Direct materials Direct labor Manufacturing overhead applied Total job cost $ 1,922,000 Job Omega Direct materials Direct labor Manufacturing overhead applied $ 287,900 405,900 233,700 $927,500 Total job cost Required 1. Calculate the plantwide predetermined overhead rate 2. Complete the job cost sheet for Job AlphaComplete this question by entering your answers in the tabs below Required 1 Required 2 Calculate the plantwide predetermined overhead rate. (Round your answer to 2 decimal places.) Plantwide predetermined overhead rate per DLH Required 1 Required 2 >
Answer #1

Answer

Job hours used in job omega is 24600 ( 405900/16.50)

Predetermined overhead =9.5 (233700/24600)

Labor cost = 61700*16.50 = 1018050

Manufacturing overhead = 61700*9.5 = 586150

Direct materials = total cost - labor cost - manufacturing overhead

Direct material = 1992000-1018050-586150 = 387800

Job alpha
Direct material 387800
Direct labor 1018050
Manufacturing oh 586150
Total job cost 1992000